# Bank–Laine functions with preassigned number of zeros

### v2

@@ 1. Introduction
To complete the proof, we determine the order of the entire function $A(z)$. Since $E$ is of finite order, it follows by the relation (1.8) together with (2.18) and the the lemma on the logarithmic derivative that
@@
Thus the first main theorem implies that $\rho(A)\leq 1/\text{Re}(\mu)$. On the other hand, we note that the function $U$ defined in (2.10) is meromorphic in $D^{+}$ and, in particular, is analytic on $\Gamma$ and $\Gamma^{\prime}$ respectively. By taking the derivative of $F$, we get
$\begin{split}\frac{F^{\prime}}{F}=\left(\frac{g^{\prime}_{m_{2},n_{2}}}{g_{m_{2},n_{2}}}\circ h\circ\psi^{-1}\right)(h^{\prime}\circ\psi^{-1})(\psi^{-1})^{\prime}.\end{split}$
We choose a curve in $\psi^{-1}(G^{+})$, say $\Gamma^{\prime\prime}$, that is close to the curve $\Gamma^{\prime}$. Then, by the definition of $g_{m,n}$ in (1.5) and the relation in (2.11), we see that $E(z)$ satisfies $1/E(z)\sim(1+o(1))e^{((1+o(1))z)^{1/\mu}}((1/\mu+o(1))z)^{(1-\mu)/\mu}$ as $z\to\infty$ on the curve $\Gamma^{\prime\prime}$. It follows by the relation in (1.8) that $\log|A(z)|\geq d|z|^{1/\text{Re}(\mu)}$ for some positive constant $d$ and all $z$ on the curve $\Gamma^{\prime\prime}$. Thus we also have $\rho(A)\geq 1/\text{Re}(\mu)$. Hence $\rho(A)=1/\text{Re}(\mu)$. This completes the proof of Theorem 1.3.
@@ 3.1. Preliminary lemmas
Let $\lambda\in[0,1)$. Then there is a sequence $(m_{k})$ of integers 0 and 1 only, with $m_{k}=0$ for $k=1,2,3$, such that the function $h:[0,\infty)\to[0,\infty)$ which satisfies $h(0)=0$ and which is linear in the intervals $Y_{k}:[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ there satisfies
@@
The statement in [18, Lemma 2.2] contains an error, so here we present a complete proof for the lemma. Moreover, we will construct the sequence $(m_{k})$ including infinitely many elements $1$ for the case $\lambda 0$.
@@
Let $\lambda_{1},\lambda_{2}\in[0,1]$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. By Lemma 3.1, for the number $\lambda_{1}\in[0,1)$, there is a sequence $(m_{k})$ of 00 and $1$ only and a function $h_{1}:[0,\infty)\to[0,\infty)$ which satisfies $h_{1}(0)=0$ and which is linear in the intervals $Y_{k}:[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ there satisfies
@@
Also, for the number $\lambda_{2}\in[0,1)$, there is a sequence $(n_{k})$ of 00 and $1$ only and a function $h_{2}:[0,\infty)\to[0,\infty)$ which satisfies $h_{2}(0)=0$ and which is linear in the intervals $Y_{k}:[2\pi(k-1),2\pi k]$ and has slopes $2n_{k}$ there satisfies
@@
For each pari $(m_{k},2n_{k})$, we denote $N_{k}=m_{k}+2n_{k}+1$. Then, for the two sequences $(m_{k})$ and $(n_{k})$ chosen, a simple calculation shows that the function $H:[0,\infty)\to[0,\infty)$ which satisfies $H(0)=0$ and which is linear in the intervals $Y_{k}:[2\pi(k-1),2\pi k]$ and has slopes $N_{k}$ there satisfies
@@
where $\omega:[0,\infty)\to[0,\infty)$ is a function satisfying $\omega(x)\sim 2(\log x)^{2}$ as $x\to\infty$ when $\lambda_{1}=\lambda_{2}=0$ or $\omega(x)\sim 2x^{\lambda_{2}}$ as $x\to\infty$ when $\lambda_{1}<\lambda_{2}$ or $\omega(x)\sim 3x^{\lambda_{2}}$ as $x\to\infty$ when $\lambda_{1}=\lambda_{2}>0$. It is easy to see that the distance between the curve determined by the function $y=H(x)$ and the curve determined by the function $y=x$ tends to infinity as $x\to\infty$. This makes it impossible to find a quasiconformal mapping as the one $Q$ in the plane to glue as in [18, Subsection 2.2].
We shall expand the intervals $Y_{k}:[2\pi(k-1),2\pi k]$ suitably to obtain a function $g:[0,\infty)\to[0,\infty)$ which satisfies $g(0)=0$ and which is linear in the intervals $Y_{k}:[2\pi(k-1),2\pi k]$ such that
$\begin{split}H(g(x))=x,\quad x\to\infty.\end{split}$
Such a function $g$ exists and, by the estimates in (3.8), we see that $g(x)=x(1+o(1))$ as $x\to\infty$. To determine one such $g$ explicitly, we denote $\mathcal{N}_{k}=\sum_{j=1}^{k}N_{j}$ and modify the previous functions $h_{1}$, $h_{2}$ and $H$ in the intervals
$\begin{split}\mathcal{Y}_{k}:[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]\end{split}$
to be linear there and has slops $m_{k}/N_{k}$, $2n_{k}/N_{k}$ and $1$, respectively. Then, for the constant $\lambda_{1}\in[0,1)$ and the sequence $(m_{k})$ before, there is a function $\tilde{h}_{1}:[0,\infty)\to[0,\infty)$ which satisfies $\tilde{h}_{1}(0)=0$ and which is linear in the intervals $\mathcal{Y}_{k}:[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ and has slopes $m_{k}/N_{k}$ there satisfies
$\begin{split}\tilde{h}_{1}(x)=h_{1}(x)+O(1),\quad x\to\infty.\end{split}$
@@
(3.9) $\begin{split}c_{2}e^{(2\pi i)^{1/2}}\leq 2\pi k_{i}\leq c_{1}e^{(2\pi i)^{1/2}},\quad i\to\infty\end{split}$
@@
(3.10) $\begin{split}c_{4}(2\pi i)^{1/\lambda_{1}}\leq 2\pi k_{i}\leq c_{3}(2\pi i)^{1/\lambda_{1}},\quad i\to\infty\end{split}$
for two positive constants $c_{3},c_{4}$. Also, for the constant $\lambda_{2}\in[0,1)$ and the sequence $(n_{k})$ before, there is a function $\tilde{h}_{2}:[0,\infty)\to[0,\infty)$ which satisfies $\tilde{h}_{2}(0)=0$ and which is linear in the intervals $\mathcal{Y}_{k}:[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ and has slopes $2n_{k}/N_{k}$ there satisfies
$\begin{split}\tilde{h}_{2}(x)=h_{2}(x)+O(1),\quad x\to\infty.\end{split}$
@@
(3.11) $\begin{split}c_{6}e^{(2\pi j)^{1/2}}\leq 2\pi k_{j}\leq c_{5}e^{(2\pi j)^{1/2}},\quad j\to\infty\end{split}$
@@
(3.12) $\begin{split}c_{8}(2\pi j)^{1/\lambda_{2}}\leq 2\pi k_{j}\leq c_{7}(2\pi j)^{1/\lambda_{2}},\quad j\to\infty\end{split}$
@@
The proof for large $N$ has been given in [18, Lemma 2.4]. We write $s_{m,n}=\log N+r$. Then equation (1.5) yields
@@ 3.2. Definition of a quasiregular map
Now we begin to define a quasiregular map by gluing the functions $g_{m,n}$ defined in (1.5) for two suitably chosen sequences $(m_{k})$ and $(n_{k})$. By modifying the functions $g_{m,n}$ suitably, we obtain closely related functions $u_{m,n}$ and $v_{m,n}$ and then glue restrictions of these maps to strips along horizontal lines to obtain quasiregular maps $U$ and $V$ which are defined in the right and left half-planes respectively. In particular, we have $U(iy)=V(iy)$ on the imaginary axis, resulting in a quasiregular map $G$ in the plane. The maps $U$, $V$ and $G$ will commute with complex conjugation, so it will be enough to define them in the upper half-plane.
@@
(3.20) $\begin{split}\tilde{g}_{m_{k},n_{k}}(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(x+i\frac{y-2\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x+iy-2i\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\leq 0.\end{array}\right.\end{split}$
@@
(3.21) $\begin{split}\mathbf{\Pi}_{k}^{+}=\{x+iy:x>0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}.\end{split}$
@@
which is continuous along the horizontal lines $\mathcal{L}^{+}_{k}$ in the right half-plane and $\mathcal{L}^{-}_{k}$ in the left half-plane for any integer $k$ respectively. In the section below, we will show that the map $G$ defined in (3.25) is indeed quasiregular in the plane and also satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem [13, § V.6].
@@ 3.3. Estimation of the dilatation
By the explicit expressions of $\tilde{g}_{m_{k},n_{k}}$ in (3.20), the arguments for estimating the complex dilation of $G$ will be similar in either of the three cases (I), (II) and (III). Below we shall first consider the first two cases (I) and (II) together.
@@
In order to estimate $K_{U}(z)-1$ for $z\in\mathbf{\Pi}_{k}^{+}$ defined in (3.21), we recall from Lemma 2.1 that
@@
The last two equations imply that the functions $r(x)$ defined by (3.29) satisfies $r(x)=O(1)$ for $z\in\mathbf{\Pi}_{k}^{+}$. Moreover, we note that the affine mapping $\chi_{m_{k},n_{k}}:H^{+}\to H^{+}$ defined in (3.18) satisfies $\chi_{m_{k},n_{k}}(z)=x/l+iy/N_{k}$ and so $|\mu_{\chi_{m_{k},n_{k}}}|=(N_{k}-l)/(N_{k}+l)<1$. Thus $U$ is quasiregular in the right half-plane $H^{+}$ and
@@
$\begin{split}\int_{S_{1,k_{i}}}\frac{dxdy}{x^{2}+y^{2}}\leq\frac{\pi}{A(2\pi i)^{1/\gamma}}\end{split}$
for $i\geq 1$ and some positive constant $A$. Thus
@@
for all $i\geq 1$ and some constant $A_{1}$. Similarly, we also have
@@
for all $i\geq 1$ and some constant $A_{2}$. Moreover, together with the the estimates in (3.11) and (3.12), we can also deduce that
@@
for all $j\geq 1$ and some constant $A_{3}$, and that
@@
for all $j\geq 1$ and some constant $A_{4}$. Note that $U$ is meromorphic in $H^{+}$ apart from the half-strips $S_{1,k_{i}-1}$, $S_{1,k_{i}}$, $S_{1,k_{i}+1}$, $S_{1,k_{j}-1}$, $S_{1,k_{j}}$ and $S_{1,k_{j}+1}$. Then the last four inequalities (3.32)-(3.35) imply that
@@
The last two inequalities imply that the function $r(x)$ defined in (3.37) satisfies $r(x)=O(1)$ for $z\in\mathbf{\Pi}_{k}^{-}$ defined in (3.23).
@@
For the case (III), since the two sequences $(m_{k})$ and $(n_{k})$ are chosen in the way that $m_{k}=n_{k}=1$ for all $k\geq 4$, we have $N_{k}=4$ for all $k\geq 4$. Thus we only need to estimate the complex dilatation of $U$ and $V$ in the stripes $\mathbf{\Pi}_{k}^{+}$ and $\mathbf{\Pi}_{k}^{-}$ for the case $k=1,2,3$ to conclude that the inequality in (3.39) also holds.
Therefore, in either of the three cases (I), (II) and (III), the function $G$ defined in (3.25) satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem [13, § V.6]. This theorem, together with the existence theorem for quasiconformal mappings [13, § V.1], yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
(3.40) $\begin{split}G(z)=F(\tau(z))\ \ \text{and}\ \ \tau(z)\sim z\ \ \text{as}\ z\to\infty.\end{split}$
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A(z)$.

### v3

@@ 1. Introduction
Finally, we determine the order of the entire function $A$. Since $E$ is of finite order, it follows by the relation (1.8) together with (2.18) and the lemma on the logarithmic derivative that
@@
Thus the first main theorem implies that $\rho(A)\leq 1/\text{Re}(\mu)$. On the other hand, we note that the function $U$ defined in (2.10) is meromorphic in $G^{+}$ and, in particular, is analytic on $\Gamma$ and $\Gamma^{\prime}$ respectively. By taking the derivative of $F$ in $\tau(\hat{G})$, we get
(2.19) $\begin{split}\frac{F^{\prime}}{F}=\left(\frac{g^{\prime}_{m_{2},n_{2}}}{g_{m_{2},n_{2}}}\circ h\circ\tau^{-1}\right)(h^{\prime}\circ\tau^{-1})(\tau^{-1})^{\prime}.\end{split}$
We may choose a curve in $\tau(\hat{G})$, say $\Gamma^{\prime\prime}$, so that $\tau^{-1}(\Gamma^{\prime\prime})$ is close to the curve $\Gamma^{\prime}$. For a point $z\in\tau^{-1}(\Gamma^{\prime\prime})$, we have the Cauchy formula
[…]
with a circle $C_{z}$ centered at $z$. Choosing the radius $\beta(z)$ of this circle to be finite, say $\beta(z)=1$, we have $|\tau^{\prime}(z)|\leq 2|z|$ when $|z|$ is large. Since $E=F/F^{\prime}$, then by the definition of $g_{m,n}$ in (1.5) as well as the relations in (2.11) and in (2.19), we see that $1/E$ satisfies
(2.20) $\begin{split}\frac{1}{|E(z)|}\geq\left(d_{1}e^{\frac{1}{2}|z|^{1/\text{Re}(\mu)}}\right)\left(\frac{d_{2}|z|^{1/\text{Re}(\mu)-1}}{|\mu|}\right)\frac{1}{2|z|}\end{split}$
for two suitably positive constants $d_{1},d_{2}$ and all $z\in\Gamma^{\prime\prime}$ such that $|z|$ is large. Further, since $E$ is of finite order, by an estimate in [7, Corollary 2] we may suppose that, for a small $\varepsilon>0$, there is an infinite sequence $(z_{n})$ such that $z_{n}\in\Gamma^{\prime\prime}$ and $|E^{\prime}(z_{n})/E(z_{n})|\leq|z_{n}|^{1/\text{Re}(\mu)-1+\varepsilon}$ and $|E^{\prime\prime}(z_{n})/E(z_{n})|\leq|z_{n}|^{2(1/\text{Re}(\mu)-1+\varepsilon)}$ for all $z_{n}$ such that $|z_{n}|$ is large. Together with these two estimates as well as the inequality in (2.20), the relation in (1.8) implies that $\log|A(z_{n})|\geq d|z_{n}|^{1/\text{Re}(\mu)}$ for some positive constant $d$ and all $z_{n}\in\Gamma^{\prime\prime}$ such that $|z_{n}|$ is large, implying that $\rho(A)\geq 1/\text{Re}(\mu)$. Hence $\rho(A)=1/\text{Re}(\mu)$. This completes the proof of Theorem 1.3.
@@ 3.1. Preliminary lemmas
Let $\lambda\in[0,1)$. Then there is a sequence $(m_{k})$ of integers 0 and 1 only, with $m_{k}=0$ for $k=1,2,3$, such that the function $h:[0,\infty)\to[0,\infty)$ which satisfies $h(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ there satisfies
@@
The statement in [19, Lemma 2.2] contains an error, so here we present a complete proof for the lemma. Moreover, we will construct the sequence $(m_{k})$ including infinitely many elements $1$ for the case $\lambda=0$.
@@
Let $\lambda_{1},\lambda_{2}\in[0,1)$ be two numbers such that $\lambda_{1}\leq\lambda_{2}$. By Lemma 3.1, for the number $\lambda_{1}\in[0,1)$, there is a sequence $(m_{k})$ of 00 and $1$ only and a function $h_{1}:[0,\infty)\to[0,\infty)$ which satisfies $h_{1}(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $m_{k}$ there satisfies
@@
For the number $\lambda_{2}\in[0,1)$, there is a sequence $(n_{k})$ of 00 and $1$ only and a function $h_{2}:[0,\infty)\to[0,\infty)$ which satisfies $h_{2}(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $2n_{k}$ there satisfies
@@
For each pair $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$, we denote $N_{k}=m_{k}+2n_{k}+1$. Then, for the two sequences $(m_{k})$ and $(n_{k})$ chosen, a simple calculation shows that the function $H:[0,\infty)\to[0,\infty)$ which satisfies $H(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ and has slopes $N_{k}$ there satisfies
@@
where $\omega:[0,\infty)\to[0,\infty)$ is a function satisfying $\omega(x)\sim 2(\log x)^{2}$ as $x\to\infty$ when $\lambda_{1}=\lambda_{2}=0$ or $\omega(x)\sim 2x^{\lambda_{2}}$ as $x\to\infty$ when $\lambda_{1}<\lambda_{2}$ or $\omega(x)\sim 3x^{\lambda_{2}}$ as $x\to\infty$ when $\lambda_{1}=\lambda_{2}>0$. It is easy to see that the distance between the curve determined by the function $y=H(x)$ and the curve determined by the function $y=x$ tends to infinity as $x\to\infty$. This makes it impossible to find a quasiconformal mapping as the one $Q$ in the plane to glue as in [19, Subsection 2.2].
We shall expand the intervals $[2\pi(k-1),2\pi k]$ suitably to obtain a function $g:[0,\infty)\to[0,\infty)$ which satisfies $g(0)=0$ and which is linear in the intervals $[2\pi(k-1),2\pi k]$ such that
$\begin{split}H(g(x))=x.\end{split}$
Such a function $g$ exists and, by the estimates in (3.8), we see that $g(x)=x(1+o(1))$ as $x\to\infty$. To determine one such $g$ explicitly, we denote $\mathcal{N}_{k}=\sum_{j=1}^{k}N_{j}$ and modify the previous functions $h_{1}$, $h_{2}$ and $H$ in the intervals $[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ to be linear there and has slops $m_{k}/N_{k}$, $2n_{k}/N_{k}$ and $1$, respectively. Then, for the constant $\lambda_{1}\in[0,1)$ and the sequence $(m_{k})$ before, there is a function $H_{1}:[0,\infty)\to[0,\infty)$ which satisfies $H_{1}(0)=0$ and which is linear in the intervals $[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ and has slopes $m_{k}/N_{k}$ there satisfies
$\begin{split}H_{1}(x)=h_{1}(x)+O(1),\quad x\to\infty.\end{split}$
@@
(3.9) $\begin{split}c_{2}e^{(2\pi i)^{1/2}}\leq 2\pi k_{i}\leq c_{1}e^{(2\pi i)^{1/2}}\end{split}$
@@
(3.10) $\begin{split}c_{4}(2\pi i)^{1/\lambda_{1}}\leq 2\pi k_{i}\leq c_{3}(2\pi i)^{1/\lambda_{1}}\end{split}$
for two positive constants $c_{3},c_{4}$. Also, for the constant $\lambda_{2}\in[0,1)$ and the sequence $(n_{k})$ before, there is a function $H_{2}:[0,\infty)\to[0,\infty)$ which satisfies $H_{2}(0)=0$ and which is linear in the intervals $[2\pi\mathcal{N}_{k-1},2\pi\mathcal{N}_{k}]$ and has slopes $2n_{k}/N_{k}$ there satisfies
$\begin{split}H_{2}(x)=h_{2}(x)+O(1),\quad x\to\infty.\end{split}$
@@
(3.11) $\begin{split}c_{6}e^{(2\pi j)^{1/2}}\leq 2\pi k_{j}\leq c_{5}e^{(2\pi j)^{1/2}}\end{split}$
@@
(3.12) $\begin{split}c_{8}(2\pi j)^{1/\lambda_{2}}\leq 2\pi k_{j}\leq c_{7}(2\pi j)^{1/\lambda_{2}}\end{split}$
@@
The proof for large $N$ has been given in [19, Lemma 2.4]. We write $s_{m,n}=\log N+r$. Then equation (1.5) yields
@@ 3.2. Definition of a quasiregular map
Now we begin to define a quasiregular map by gluing the functions $g_{m,n}$ defined in (1.5) for two suitably chosen sequences $(m_{k})$ and $(n_{k})$. By modifying the functions $g_{m,n}$ suitably, we obtain closely related functions $u_{m,n}$ and $v_{m,n}$ and then glue restrictions of these maps to strips along horizontal lines to obtain quasiregular maps $U$ and $V$ which are defined in the right and left half-planes respectively. In particular, we have $U(iy)=V(iy)$ on the imaginary axis, resulting in a quasiregular map $G$ in the plane. The maps $U$, $V$ and $G$ will commute with the complex conjugation, so it will be enough to define them in the upper half-plane.
@@
(3.20) $\begin{split}\tilde{g}_{m_{k},n_{k}}(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(\frac{x}{l}+i\frac{y-2\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x+iy-2i\pi\mathcal{N}_{k-1}}{N_{k}}\right),&x\leq 0.\end{array}\right.\end{split}$
@@
(3.21) $\begin{split}\hat{\mathbf{\Pi}}_{k}^{+}=\{x+iy:x>0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}.\end{split}$
@@
which is continuous along the horizontal lines $\mathcal{L}^{+}_{k}$ in the right half-plane and $\mathcal{L}^{-}_{k}$ in the left half-plane for any integer $k$ respectively. In the section below, we will show that the map $G$ defined in (3.25) is indeed quasiregular in the plane and also satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem [14, § V.6].
@@ 3.3. Estimation of the dilatation
By the explicit expressions of $\tilde{g}_{m_{k},n_{k}}$ in (3.20), the arguments for estimating the dilation of $G$ will be similar in either of the three cases (I), (II) and (III). Below we shall first consider the first case (I).
@@
In order to estimate $K_{U}(z)-1$ for $z\in\hat{\mathbf{\Pi}}_{k}^{+}$ defined in (3.21), we recall from Lemma 2.1 that
@@
The last two equations imply that the functions $r(x)$ defined by (3.29) satisfies $r(x)=O(1)$ for $z\in\hat{\mathbf{\Pi}}_{k}^{+}$. Moreover, we note that the affine mapping $\chi_{m_{k},n_{k}}:H^{+}\to H^{+}$ defined in (3.18) satisfies $\chi_{m_{k},n_{k}}(z)=x/l+iy/N_{k}$ and so $|\mu_{\chi_{m_{k},n_{k}}}|=(N_{k}-l)/(N_{k}+l)<1$. Thus $U$ is quasiregular in the right half-plane $H^{+}$ and
@@
$\begin{split}\int_{S_{1,k_{i}}}\frac{dxdy}{x^{2}+y^{2}}\leq\frac{\pi}{c(2\pi i)^{1/\gamma}}\end{split}$
for $i\geq 1$ and some positive constant $c$. Thus
@@
for all $i\geq 1$ and some constant $c_{1}$. Similarly, we also have
@@
for all $i\geq 1$ and some constant $c_{2}$. Moreover, together with the estimates in (3.11) and (3.12), we can also deduce that
@@
for all $j\geq 1$ and some constant $c_{3}$, and that
@@
for all $j\geq 1$ and some constant $c_{4}$. Note that $U$ is meromorphic in $H^{+}$ apart from the half-strips $S_{1,k_{i}-1}$, $S_{1,k_{i}}$, $S_{1,k_{j}-1}$ and $S_{1,k_{j}}$. Then the last four inequalities (3.32)-(3.35) imply that
@@
The last two inequalities imply that the function $r(x)$ defined in (3.37) satisfies $r(x)=O(1)$ for $z\in\hat{\mathbf{\Pi}}_{k}^{-}$ defined in (3.23).
@@
For the case (II), since $l=3$ and since the sequence $(n_{k})$ are chosen in the way that $n_{k}=1$ for all $k\geq 4$, we also have we have $N_{k}\leq 4$ for all $k\geq 4$ and we only need to estimate the complex dilatation of $U$ in the stripes $S_{1,k_{i}-1}$ and $S_{1,k_{i}}$ of $\hat{\mathbf{\Pi}}_{k}^{+}$ and the dilation of $V$ in the stirps $T_{1,k_{i}-1}$ and $T_{1,k_{i}}$ of $\hat{\mathbf{\Pi}}_{k}^{-}$ as well as the complex dilatation of $U$ in the stripes $\hat{\mathbf{\Pi}}_{k}^{+}$ for the case $k=1,2,3$ to conclude that the inequality in (3.39) also holds.
For the case (III), since $l=4$ and since the two sequences $(m_{k})$ and $(n_{k})$ are chosen in the way that $m_{k}=n_{k}=1$ for all $k\geq 4$, we have $N_{k}=4$ for all $k\geq 4$. Thus we only need to estimate the complex dilatation of $U$ in the stripes $\hat{\mathbf{\Pi}}_{k}^{+}$ for the case $k=1,2,3$ to conclude that the inequality in (3.39) also holds.
Therefore, in either of the three cases (I), (II) and (III), the function $G$ defined in (3.25) satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem. This theorem, together with the existence theorem for quasiconformal mappings, yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.